Circular No. 8908

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 Y2 (McNAUGHT) R. H. McNaught reports his discovery of a comet (as tabulated below)

on CCD images taken with the 0.5-m Uppsala Schmidt telescope, where the object appears marginally, but consistently, diffuse with a coma diameter of ∼ 10′′ in good seeing. Three stacked 90-s R images taken by D. M. Burton, Australian National University, with the ANU 1.0-m f/8 reflector at Siding Spring on 2008 Jan. 1.69 show a slightly diffuse comet (slightly involved with a star) with a 3′′ tail in p.a. 20^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, C. Jacques and E. Pimentel write that their astrometric CCD images from Jan. 1.14 (taken with a 0.30-m f/3 reflector at Belo Horizonte, Brazil) show a central condensation of size 10′′.

2007 UT  α_{2000}  δ_{2000}  Mag.
Dec. 31.66576  6^{h}45^{m}01.^{s}58  −50^{o}45′51′′.5  19.0

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-A6.

T = 2008 May 18.189 TT  ω = 264.^{o}435
Ω = 303.854 }2000.0
q = 4.13214 AU  i = 98.702

SUPERNOVA 2007uy IN NGC 2770 S. Nakano, Sumoto, Japan, reports the discovery of an apparent su-

pernova (mag 17.2) by Yoji Hirose (Tsutsumi, Chigasaki, Kanagawa-ken, Japan) on ten unfiltered CCD frames taken around 2007 Dec. 31.669 UT using a 0.35-m f/6.8 Schmidt-Cassegrain telescope (limiting mag 18.0); Nakano has measured mag 17.0 and the following position for the new object from a jpeg image supplied by Hirose: α = 9^{h}09^{m}35^{s}.40, δ = +33^{o}07′09′′.9 (equinox 2000.0), which is 20′′.6 east and 15′′.5 south of the center of NGC 2770. Nothing is visible at this position on red Digitized Sky Survey images from 1998 and 2000, though there are apparently numerous H II regions nearby. Additional unfiltered CCD magnitudes for 2007uy: 2007 Mar. 7 and Dec. 18, [18 (Hirose); Dec. 31.786, 17.3 (K. Kadota, Ageo, Japan, 0.25m f/5 reflector; position end figures 35^{s}.27, 08′′.9); 2008 Jan. 1.595, 17.0 (Hirose); 1.651, 16.9 (K. Itagaki, Takanezawa station, Tochigi-ken, Japan, 0.30-reflector; position end figures 35^{s}.35, 08′′.9); 1.885, 16.8 (P. Corelli, Pagnacco, Italy, 0.45-m f/4.5 reflector; position end figures 35^{s}.17, 08′′.9).

2008 January 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green